# VERUS-LM: a Versatile Framework for Combining LLMs with Symbolic Reasoning

### v1

@@ Abstract
A recent approach to neurosymbolic reasoning is to explicitly combine the strengths of large language models (LLMs) and symbolic solvers to tackle complex reasoning tasks. However, current approaches face significant limitations, including poor generalizability due to task-specific prompts, inefficiencies caused by the lack of separation between knowledge and queries, and restricted inferential capabilities. These shortcomings hinder their scalability and applicability across diverse domains. In this paper, we introduce VERUS-LM, a novel framework designed to address these challenges. VERUS-LM employs a generic prompting mechanism, clearly separates domain knowledge from queries, and supports a wide range of different logical reasoning tasks. This framework enhances adaptability, reduces computational cost, and allows for richer forms of reasoning, such as optimization and constraint satisfaction. We show that our approach succeeds in diverse reasoning on a novel dataset, markedly outperforming LLMs. Additionally, our system achieves competitive results on common reasoning benchmarks when compared to other state-of-the-art approaches, and significantly surpasses them on the difficult AR-LSAT dataset. By pushing the boundaries of hybrid reasoning, VERUS-LM represents a significant step towards more versatile neurosymbolic AI systems. [^1] All code and datasets required to reproduce the results in this text are available online: https://gitlab.com/EAVISE/bca/verus-lm
@@ 1 Introduction
Logical reasoning is an essential aspect of problem-solving, decision-making, and critical thinking, making it an important goal in Artificial Intelligence. While recent large language models (LLMs) such as GPT-4 Achiam et al. (2023) have shown an impressive leap in reasoning capabilities through techniques like chain-of-thought prompting, they often lack the ability to guarantee faithful and transparent reasoning. Indeed, at a fundamental level, LLMs operate as black-box probabilistic models, which makes it difficulty to ensure accurate and coherent reasoning steps.
Symbolic inference engines, on the other hand, can derive conclusions that are provably correct and typically also explainable. However, these systems struggle to interpret ambiguous natural language input. This creates an opportunity for neurosymbolic methods to combine the strengths of LLMs with symbolic reasoning systems. Interestingly, this method loosely parallels Kahneman’s theory on human cognition, in which reasoning is divided into fast and unconscious thinking (System 1) and slow and methodical thinking (System 2) Kahneman (2011). These approaches have already demonstrated impressive improvements on common reasoning datasets Pan et al. (2023); Ishay et al. (2023); Olausson et al. (2023).
Despite these promising results, current state-of-the-art approaches have several limitations that hinder their applicability across diverse domains. One significant drawback is their heavy reliance on task-specific prompts to generate symbolic representations, via in-context learning. This makes these approaches less general and harder to adapt to new domains, since this may require manual prompt engineering. In real-world scenarios, where problems are highly variable, task-specific prompts can cause a severe bottleneck. Furthermore, they typically try to reduce every user query to the same logical reasoning task (or to one of only a very limited set of reasoning tasks). In many domains, users have widely different kinds of queries, which correspond naturally to equally wide range of logical reasoning tasks (e.g., such as optimization, generating solutions, deriving consequences, and generating explanations). Therefore, this lack of versatility limits the effectiveness of the state-of-the-art tools in complex domains where multiple forms of reasoning are required.
In this paper, we propose VERUS-LM, an enhanced neurosymbolic framework that addresses these limitations. It features a generic prompting pipeline, a semantic refinement step, a clear separation between declarative knowledge from questions, and support for a wide range of logical reasoning tasks. In this way, it is able to handle complex and dynamic environments in a robust and effective manner.
@@ 2 Related Work
Improving the reasoning capabilities of LLMs has recently become a topic of great interest. For the purpose of this paper, however, we will not discuss works on improving LLM reasoning with in-context prompting or by fine-tuning, but refer to Huang and Chang (2023); Plaat et al. (2024) for more information on those techniques. Instead, we will focus on the neurosymbolic approach in which an LLM is coupled with an explicit symbolic reasoning engine.
@@
Current state-of-the-art systems all largely work in the same way. To solve a given reasoning problem, which typically consists of a description of some domain and a query that needs to be answered, they first let an LLM generate a formal representation of the problem and then pass this on to a reasoning engine to solve. In this way, they combine the benefits of both kinds of AI systems: the LLM offers a user-friendly natural language interface, while the reasoning engine ensures correct reasoning. This offers a promising path towards improving the reasoning skills of LLM-based systems. We will now briefly go over some of these systems and touch on how they differ.
@@
In Ishay et al. (2023), the authors present a pipeline to automatically generate Answer Set Programming (ASP) representations Brewka et al. (2011). Their method uses the LLM three times: first to extract relevant constants, next to generate predicates, and finally to generate the ASP rules. Using a logic puzzle dataset, they demonstrate a 71% increase in accuracy compared to a baseline GPT-4. The same authors also rely on ASP in the [LLM]+ASP framework Yang et al. (2023), where given a problem description, the LLM extracts a set of atomic facts. These facts, coupled with pre-defined, hand-crafted “knowledge modules” containing domain-specific ASP rules, are fed to an ASP solver to find a solution. They validate their approach on four “story-based” datasets, on which it outperforms the state-of-the-art.
@@
In the LINC system Olausson et al. (2023), an LLM is used to generate statements in First Order Logic (FOL), which are given to a FOL reasoner. To mitigate formalization errors, they generate $K$ formalizations and make use of $K$-way majority voting to decide on the correct response. They demonstrate that their approach is competitive on the FOLIO dataset Han et al. (2022), and achieves remarkable improvements on the ProofWriter dataset Tafjord et al. (2020).
@@
The final neurosymbolic system we discuss, Logic-LM Pan et al. (2023), has two distinctive features. First, instead of relying on a single formalization language, their system supports logic programs, FOL, constraint satisfaction problems and SAT encodings. In this way, the system is able to handle a broader range of problem types. Second, they introduce a self-refinement stage, in which the LLM iteratively attempts to fix any syntax errors based on feedback from the solver. They validate their approach on five datasets, on which they generally outperform GPT-4 using chain-of-thought as baseline.
@@
It is worth noting that most of the aforementioned methods rely heavily on dataset-specific examples, utilized for in-context learning, to guide the LLMs. Furthermore, they only support a limited number of logical reasoning tasks, which limits their generalization capabilities to new datasets. This is especially apparent on the AR-LSAT dataset Zhong et al. (2021), featuring logical reasoning questions from the Law School Admission Test, which is high in diversity and thus requires broad reasoning capabilities. For instance, given knowledge on the types of CDs sold by a store, the questions can range from “Which CDs are on sale?” to “Given $\phi$, which statements are true?” and “What are the minimum CDs required to be on sale if $\phi$?”, for some additional statement $\phi$. Due to this diversity, Logic-LM only achieves a 43.04% accuracy on this dataset, though still an improvement over baseline GPT-4.
@@
While it is strictly speaking not a neurosymbolic approach, we also briefly discuss the SymbCoT framework Xu et al. (2024) for the sake of comparison. Like the aforementioned works, SymbCoT internally translates a problem into FOL using in-context learning. However, instead of using a separate reasoning engine, the LLM itself then performs logical reasoning on the generated statements in a step-by-step way similar to Chain-of-Thought, instructed by specific prompts that try to mimic the kinds of logical reasoning found in logical solvers. By not using a separate reasoning engine, the approach becomes more robust to syntax errors. Interestingly, this approach consistently outperforms standard Chain-of-Thought and Logic-LM on the PrOntoQA Saparov and He (2022), FOLIO and ProofWriter datasets.
@@ 3.1 Requirements
As discussed in the previous section, current state-of-the-art approaches typically focus on a single reasoning tasks. However, more realistic use cases will likely require a more versatile approach. Imagine, for instance, a chat bot acting as a digital assistant in the field of car insurance. A person interacting with such an AI system will not only ask questions like “Am I eligible for an insurance policy?”, but will also ask other things such as “Why am I (in)eligible?” or “Depending on my car type, what would my insurance premium?”.
@@
    Versatile: it should support multiple forms of reasoning (e.g., verification, explanation, optimisation, etc.)
@@ 3.2 Reasoning Engine
While the aforementioned requirements pertain to the general design of our system, they also constrain our selection of reasoning engine. To serve as the main reasoning core of VERUS-LM, we have selected the IDP-Z3 system Carbonnelle et al. (2023). IDP-Z3 is a reasoning engine for FO($\cdot$), a rich extension of classical First-Order Logic with useful features such as types, aggregates, (inductive) definitions, and more. It explicitly supports the Knowledge Base Paradigm Denecker and Vennekens (2008), in which the same domain knowledge can be used for many different reasoning tasks.
We will briefly illustrate FO($\cdot$) and the forms of reasoning supported by IDP-Z3 through a small example, referring to Carbonnelle et al. (2023) for a more extensive overview.
As in normal FOL, a vocabulary in FO($\cdot$) consists of a set of symbols. Because FO($\cdot$) is a typed logic, this includes types, in addition to predicate and function symbols. For instance, in the below example about car insurance, we use types Customer and Car. The function risk_factor maps each Car to its associated risk factor ($\in\mathbb{R}$), while the function age maps each Applicant to their age ($\in\mathbb{Z}$). A number of constants (= 0-ary functions) represent the car’s value, its type, and the insurance premium. Finally, a unary predicate applicant represents which of the customers is requesting the insurance. Again following FOL, a structure for a vocabulary provides an interpretation for some (not necessarily all) of the symbols in this vocabulary. Finally, a theory consists of a set of logical sentences. In this simple example, we have two sentences: one states that every applicant must be an adult, and one defines the calculation of the insurance premium.
vocabulary V {
type Customer :=assign:=:= {Ann, Barbara}
type Car :=assign:=:= {Sedan, SUV, ..., Truck}
age: Applicant $\rightarrow$ Int
risk_factor: Car $\rightarrow$ Real
car_value: $\rightarrow$ Real
car_type: $\rightarrow$ Car
applicant: Customer $\rightarrow$ Bool
premium: $\rightarrow$ Real
structure S:V {
@@
theory T:V {
@@
premium() = (car_value()/100) $\times$ risk_factor(car_type()).
[…]
Given such a KB, IDP-Z3 can perform different forms of reasoning to answers different questions. For instance, IDP-Z3 can determine which customers are eligible for an insurance policy, and explain why. Given the car type and car value, it can also calculate the insurance premium. Or, it can “optimize” the premium by looking for the car type with the lowest risk factor. None of these different reasoning tasks require any modifications to the KB itself, allowing for the easy reuse of the knowledge.
These properties of IDP-Z3 make it a natural choice for our reasoning system. Additionally, FO($\cdot$) is expressive enough to model many problem domains, as supported by real-life use cases Aerts et al. (2022); Deryck et al. (2019); Vandevelde et al. (2022). Furthermore, as First-Order Logic is the subject of many articles and textbooks, the training corpora of LLMs will also contain enough information to allow them to be fluent in this formalism.
@@ 4 Knowledge Base Creation
The first phase of the VERUS-LM framework consists of transforming domain knowledge into a formal KB, inspired by LLM’s capability of translating natural language to structured formats like equations He-Yueya et al. (2023) or Python code Gao et al. (2023). This translation is performed in three steps, which we will discuss in the following sections.
@@ 4.1 Symbol Extraction
The LLM first identifies the relevant types, functions, and predicate symbols for the problem domain. We use a prompt that describes these three different kinds of symbols and how they are typically used. Important here is that we do not include examples from a specific dataset. Instead, our prompt just provides general instructions on how to identify and categorize the different concepts, and provides three general examples as an illustration. The prompt also contains the instruction to annotate each symbol with its informal language meaning, so that this information is present in the next phases of the pipeline.
@@ 4.3 Self-Refinement
Although in our experience the LLM often produces correct logical sentences, errors still occur. To correct these, we introduce a self-refinement step that uses feedback from the reasoning engine to correct erroneous logical statements. This approach is inspired by recent work on imperative programming Chen et al. (2023); Madaan et al. (2024) and the Logic-LM framework Pan et al. (2023). However, whereas these approaches only consider syntactic correctness, we introduce a novel second step to take semantic correctness into account as well, by means of a satisfiability check. These refinements are part of an iterative process, where VERUS-LM will repeat them until no errors are found or a maximum of attempts is reached.
@@
If the reasoning engine finds a syntax error, the LLM is instructed to correct its output, i.e., either the vocabulary and/or the theory it produces. To this end, the LLM is prompted with its previous erroneous output, the solver’s error message, and some examples of common syntax errors and their remedies. Detailed error descriptions are provided by FOLINT Vermeulen et al. (2022), IDP-Z3’s code analysis tool.
@@
Once the KB is syntactically correct, we use a satisfiability check to find additional semantic errors. It is reasonable to assume that the KB should be satisfiable, because, at this point, it contains only a formalization of the provided domain knowledge. For instance, if we were solving a planning problem, the KB would contain a description of the actions that can be executed, but we would not yet include a specific goal that the plan should achieve. If this KB is already inconsistent, then we take this as a sign that the LLM made a mistake in translating the domain knowledge and we instruct it to refine its previous output, based on an explanation of why this was unsatisfiable. IDP-Z3 can give such an explanation itself, by generating a “minimal unsatisfiable subset” of conflicting assignments and constraints Carbonnelle et al. (2023).

### v2

@@ Abstract
A recent approach to neurosymbolic reasoning is to explicitly combine the strengths of large language models (LLMs) and symbolic solvers to tackle complex reasoning tasks. However, current approaches face significant limitations, including poor generalizability due to task-specific prompts, inefficiencies caused by the lack of separation between knowledge and queries, and restricted inferential capabilities. These shortcomings hinder their scalability and applicability across diverse domains. In this paper, we introduce VERUS-LM, a novel framework designed to address these challenges. VERUS-LM employs a generic prompting mechanism, clearly separates domain knowledge from queries, and supports a wide range of different logical reasoning tasks. This framework enhances adaptability, reduces computational cost, and allows for richer forms of reasoning, such as optimization and constraint satisfaction. We show that our approach succeeds in diverse reasoning on a novel dataset, markedly outperforming LLMs. Additionally, our system achieves competitive results on common reasoning benchmarks when compared to similar state-of-the-art approaches, and significantly surpasses them on the difficult AR-LSAT dataset. By pushing the boundaries of hybrid reasoning, VERUS-LM represents a significant step towards more versatile neurosymbolic AI systems. All code and datasets required to reproduce the results in this text are available online: https://gitlab.com/EAVISE/bca/verus-lm
@@ 1 Introduction
Logical reasoning is an essential aspect of problem-solving, decision-making, and critical thinking, making it an important goal in Artificial Intelligence. While recent large language models (LLMs) such as GPT-4 [3] and Gemini [26] have shown an impressive leap in reasoning-like capabilities through techniques like chain-of-thought prompting, they often lack the ability to guarantee faithful and transparent reasoning. Indeed, at a fundamental level, LLMs operate as black-box probabilistic models, which makes it difficulty to ensure accurate and coherent reasoning steps. As a consequence, whether they are capable of “true reasoning” is still subject of much debate [5, 31, 23].
Symbolic inference engines, on the other hand, can derive conclusions that are provably correct and typically also explainable. However, these systems struggle to interpret ambiguous natural language input. This creates an opportunity for neurosymbolic methods to combine the strengths of LLMs with symbolic reasoning systems. Interestingly, this method loosely parallels Kahneman’s theory on human cognition, in which reasoning is divided into fast and unconscious thinking (System 1) and slow and methodical thinking (System 2) [16]. Such approaches have already demonstrated impressive improvements on common reasoning datasets [19, 15, 18].
Despite these promising results, current state-of-the-art approaches have several limitations that hinder their applicability across diverse domains. One significant drawback is their heavy reliance on task-specific prompts to generate symbolic representations, via in-context learning. This makes these approaches less general and harder to adapt to new domains, since this may require manual prompt engineering. In real-world scenarios, where problems are highly variable, task-specific prompts can cause a severe bottleneck. Furthermore, they typically try to reduce every user query to the same logical reasoning task (or to one of only a very limited set of reasoning tasks). In many domains, users have widely different kinds of queries, which correspond naturally to equally wide range of logical reasoning tasks (e.g., optimization, generating solutions, deriving consequences, and generating explanations). Therefore, there is a lack of versatility in the types of queries they are able to answer, limiting their effectiveness in complex domains where multiple forms of reasoning are required.
In this paper, we propose an enhanced neurosymbolic framework that addresses these limitations titles VERUS-LM. It features a generic prompting pipeline, a semantic refinement step, a clear separation between declarative knowledge from questions, and support for a wide range of logical reasoning tasks. In this way, it is able to handle complex and dynamic environments in a robust and effective manner.
@@ 2 Related Work
Improving the reasoning capabilities of LLMs has recently become a topic of great interest. For the purpose of this paper, however, we will not discuss works on improving LLM reasoning with in-context prompting or by fine-tuning, but refer to [14, 20] for more information on those techniques. Instead, we will focus on the neurosymbolic approach in which an LLM is coupled with an explicit symbolic reasoning engine.
@@
Current state-of-the-art systems all largely work in the same way. Given a reasoning problem, which typically consists of a description of some domain and a query that needs to be answered, they first let an LLM generate a formal representation of the problem and then pass this on to a reasoning engine to solve. In this way, they combine the benefits of both kinds of AI systems: the LLM offers a user-friendly natural language interface, while the reasoning engine ensures correct reasoning. This offers a promising path towards improving the reasoning skills of LLM-based systems. We will now briefly go over some of these systems and touch on how they differ.
@@
In [15], the authors present a pipeline to automatically generate Answer Set Programming (ASP) representations [6]. Their method uses the LLM three times: first to extract relevant constants, next to generate predicates, and finally to generate the ASP rules. Using a logic puzzle dataset, they demonstrate a 71% increase in accuracy compared to a baseline GPT-4. The same authors also rely on ASP in the [LLM]+ASP framework [33], where given a problem description, the LLM extracts a set of atomic facts. These facts, coupled with pre-defined, hand-crafted “knowledge modules” containing domain-specific ASP rules, are fed to an ASP solver to find a solution. They validate their approach on four “story-based” datasets, on which it outperforms the state-of-the-art.
@@
In the LINC system [18], an LLM is used to generate statements in First Order Logic (FOL), which are given to a FOL reasoner. To mitigate formalization errors, they generate $K$ formalizations and make use of $K$-way majority voting to decide on the correct response. They demonstrate that their approach is competitive on the FOLIO dataset [13], and achieves remarkable improvements on the ProofWriter dataset [25].
@@
Finally, though similar to the previous approaches, Logic-LM [19] has two distinctive features. First, instead of relying on a single formalization language, their system supports logic programs, FOL, constraint satisfaction problems and SAT encodings. In this way, the system is able to handle a broader range of problem types. Second, they introduce a self-refinement stage, in which the LLM iteratively attempts to fix any syntax errors based on feedback from the solver. They validate their approach on five datasets, on which they generally outperform GPT-4 using chain-of-thought as baseline.
@@
It is worth noting that most of the aforementioned methods rely heavily on dataset-specific examples, utilized for in-context learning, to guide the LLMs. Furthermore, they only support a limited number of logical reasoning tasks, which limits their generalization capabilities to new datasets. This is especially apparent on the AR-LSAT dataset [34], featuring logical reasoning questions from the Law School Admission Test, which is high in diversity and thus requires broad reasoning capabilities. For instance, given knowledge on the types of CDs sold by a store, the questions can range from “Which CDs are on sale?” to “Given $\phi$, which statements are true?” and “What are the minimum CDs required to be on sale if $\phi$?”, for some additional statement $\phi$. Though it still improves on the baseline GPT-4, Logic-LM only achieves a 43.04% accurary on this dataset due to its diversity.
@@
Though different from the others, we also briefly discuss the SymbCoT framework [32] for the sake of comparison. Like the aforementioned works, SymbCoT internally translates a problem into FOL using in-context learning. However, instead of using a separate reasoning engine, the LLM itself then performs logical reasoning on the generated statements in a step-by-step way similar to Chain-of-Thought, instructed by specific prompts that try to mimic the kinds of logical reasoning found in logical solvers. By not using a separate reasoning engine, the approach becomes more robust to syntax errors. Interestingly, this approach consistently outperforms standard Chain-of-Thought and Logic-LM on the PrOntoQA [22], FOLIO and ProofWriter datasets.
@@ 3.1 Requirements
As discussed in the previous section, current state-of-the-art approaches typically focus on a single reasoning tasks. However, more realistic use cases will likely require a more versatile approach. Imagine, for instance, a chat bot acting as a digital assistant in the field of car insurance. A person interacting with such an AI system will not only ask questions like “Am I eligible for an insurance policy?”, but will also ask other things such as “Why am I (in)eligible?”, “Depending on my car type, what would my insurance premium?” and “What elements can I change to minimize my premium?”.
@@
    Versatile: it should support multiple useful forms of reasoning (e.g., verification, explanation, optimisation, etc.)
@@ 3.2 Reasoning Engine
While the aforementioned requirements pertain to the general design of our system, they also constrain our selection of reasoning engine. To serve as the main reasoning core of VERUS-LM, we have selected the IDP-Z3 system [7]. IDP-Z3 is a reasoning engine for FO($\cdot$), a rich extension of classical First-Order Logic with useful features such as types, aggregates, (inductive) definitions, and more. It explicitly supports the Knowledge Base Paradigm [9], in which the same domain knowledge can be used for many different reasoning tasks.
We will briefly illustrate FO($\cdot$) and the forms of reasoning supported by IDP-Z3 through a small example, referring to [7] for a more extensive overview.
As in normal FOL, a vocabulary in FO($\cdot$) consists of a set of symbols. Because FO($\cdot$) is a typed logic, this includes types, in addition to predicate and function symbols. For instance, in the below example about car insurance, we use types Customer and Car. The function risk_factor maps each Car to its associated risk factor ($\in\mathbb{R}$), while the function age maps each Applicant to their age ($\in\mathbb{Z}$). A number of constants (= 0-ary functions) represent the car’s value, its type, and the insurance premium. Finally, a unary predicate applicant represents which of the customers is requesting the insurance. Again following FOL, a structure for a vocabulary provides an interpretation for some (not necessarily all) of the symbols in this vocabulary. Finally, a theory consists of a set of logical sentences. In this simple example, we have two sentences: one states that every applicant must be an adult (L18), and one defines the calculation of the insurance premium (L19).
1vocabulary V {
2 type Customer :=:= {Ann, Brit}
3 type Car :=:= {Sedan,SUV,...}
4 risk_factor: Car $\rightarrow$ Real
5 age: Applicant $\rightarrow$ Int
6 car_value: $\rightarrow$ Real
7 premium: $\rightarrow$ Real
8 car_type: $\rightarrow$ Car
9 premium: $\rightarrow$ Real
10 applicant: Customer $\rightarrow$ Bool
11}
12
13structure S:V {
14 age :=:= {Ann $\rightarrow$ 16, Brit $\rightarrow$ 32}.
@@
20 premium() = (car_value()/100) $\times$ risk_factor(car_type()).
21}
@@
Given such a KB, IDP-Z3 can perform different forms of reasoning to answers different questions. For instance, IDP-Z3 can determine which customers are eligible for an insurance policy, and explain why. Given the car type and car value, it can also calculate the insurance premium. Or, it can “optimize” the premium by looking for the car type with the lowest risk factor. Using these different reasoning tasks does not require any modifications to the KB itself, allowing for straightforward knowledge reuse.
These properties of IDP-Z3 make it a natural choice for our reasoning system. Additionally, FO($\cdot$) is expressive enough to model many problem domains, as supported by real-life use cases [4, 10, 27, 28]. Furthermore, as First-Order Logic is the subject of many articles and textbooks, the training corpora of LLMs will also contain enough information to allow them to be fluent in this formalism.
@@ 4 Knowledge Base Creation
The first phase of the VERUS-LM framework consists of transforming domain knowledge into a formal KB. This translation is performed in three steps, which we will discuss in the following sections.
@@ 4.1 Symbol Extraction
The LLM first identifies the relevant types, functions, and predicate symbols for the problem domain. We use a prompt that describes these three different kinds of symbols and how they are typically used. Important here is that we do not include examples from a specific dataset. Instead, our prompt just provides general instructions on how to identify and categorize the different concepts, and provides three general examples that do not appear in any of the datasets as an illustration. The prompt also contains the instruction to annotate each symbol with its informal language meaning, so that this information is present in the next phases of the pipeline.
@@ 4.3 Self-Refinement
Although in our experience the LLM often produces correct logical sentences, errors still occur. To correct these, we introduce a self-refinement step that uses feedback from the reasoning engine to correct erroneous logical statements. This approach is inspired by recent work on imperative programming [8, 17] and the Logic-LM framework [19]. However, whereas these approaches only consider syntactic correctness, we introduce a novel second step to take semantic correctness into account as well, by means of a satisfiability check. These refinements are part of an iterative process, where VERUS-LM will repeat them until no errors are found or a maximum of attempts is reached.
@@
If the reasoning engine finds a syntax error, the LLM is instructed to correct its output, i.e., either the vocabulary and/or the theory it produces. To this end, the LLM is prompted with its previous erroneous output, the solver’s error message, and some examples of common syntax errors and their remedies. Detailed error descriptions are provided by FOLINT [29], IDP-Z3’s code analysis tool.
@@
Once the KB is syntactically correct, we use a satisfiability check to find additional semantic errors. It is reasonable to assume that the KB should be satisfiable, because, at this point, it contains only a formalization of the provided domain knowledge. For instance, if we were solving a planning problem, the KB would contain a description of the actions that can be executed, but we would not yet include a specific goal that the plan should achieve. If this KB is already inconsistent, then we take this as a sign that the LLM made a mistake in translating the domain knowledge and we instruct it to refine its previous output, based on an explanation of why this was unsatisfiable. IDP-Z3 can give such an explanation itself, by generating a “minimal unsatisfiable subset” of conflicting assignments and constraints [7].